### 12 Mar 2024 v1

@@ 4.2 Empirical application
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.940,0.850]$ $[-0.872,0.850]$
@@
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-0.821,0.994]$ $[-0.691,0.596]$
@@
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-1,1]$
@@
$\mathbb{E}[pass_{i}|ELL_{i}=1,white_{i},FRPL_{i}]-\mathbb{E}[pass_{i}|ELL_{i}=0,white_{i},FRPL_{i}]\leq 0.$ (17)
I present estimated bounds on English exam white/non-white average pass rate gaps in Table 5. Sharp bounds on the white/non-white average pass rate gaps reported in Column 1 are again wide and either uninformative or close to uninformative. Imposing both additional monotonicity restrictions of (16) and (17) helps to narrow the bounds all parameters, reported in Column 2, but all bounds are still wide and contain 0. Thus even with additional monotonicity restrictions, it is difficult to obtain bounds that are narrow enough to be useful from realistic aggregate data alone.
Bounds without Monotonicity Monotonicity Bounds
@@
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.904,0.923]$ $[-0.823,0.923]$
@@
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-0.852,1]$ $[-0.759,0.706]$
@@
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-1,0.914]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,1]$ $[-0.759,0.706]$
@@
5 Conclusion
In this paper I present sharp bounds on individual-level parameters of interest when the only available data is aggregate data, and I develop a valid inference method relying only on the available marginal information of each covariate. In simulations and an empirical application I show that the sharp bounds are too wide to be useful with realistic aggregate data, even after imposing additional shape restrictions. This suggests that researchers hoping to use aggregate data to make individual-level claims should either obtain additional individual-level data, or impose stronger assumptions to be able to use other existing methodology. However because the sharp bounds are so wide, stronger assumptions that deliver narrower bounds will have strong implications about the underlying aggregation process, meaning the researcher should always carefully justify any imposed assumptions.
Appendix A Appendix: Proofs
A.1 Proof of Lemma 1
@@ A.2 Proof of Proposition 1
$D=\bigg{\{}\sum_{k=1}^{K}\lambda_{k}d_{k}\bigg{|}~{}0\leq d_{k}\leq 1~{}\forall k,\text{ and }\exists\,(p_{1g},\dots,p_{Kg})\in P_{g},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }d_{k}=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}~{}\forall k,~{}\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g\bigg{\}},$ (A.1)
@@
$P_{g}=\bigg{\{}(p_{1g},\dots,p_{Kg})\in[0,1]^{K}\bigg{|}~{}\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}~{}\forall\ell,k,\text{ and }\sum_{k=1}^{K}p_{kg}=1\bigg{\}}.$ (A.2)
If $\mathbb{P}[G_{i}=g],c_{kg}\in[0,1]$ for all $k$ and $g$ then because $d_{k}=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}$ we know $d_{k}\in[0,1]$ for all $k$. Then we can get rid of $d_{k}$ in (A.1) by plugging in one of the constraints like so:
$D=\bigg{\{}\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}\bigg{|}~{}\exists\,(p_{1g},\dots,p_{Kg})\in P_{g},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g\bigg{\}}.$ (A.3)
@@
This means that we can equivalently express (A.3) and (A.2) as optimization problems as follows: $D=\left[L,U\right]$, where
@@
A.3 Proof of Corollary 1

### 05 Apr 2024 v3

@@ A.3 Sharp bounds given the joint distribution
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.820,-0.580]$ $[0.531,0.860]$
@@
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-0.997,-0.814]$ $[0.795,0.981]$
@@
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-1,-0.946]$ $[0.947,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,-0.946]$ $[0.947,1]$
@@
Notes: See Section 4.1 for details about simulation exercise 1.
Table A.2: Estimated bounds under different joint covariate distributions, simulation exercise 2
Range of Lower Bounds Range of Upper Bounds
@@
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.428,-0.369]$ $[0.688,0.749]$
@@
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-1,-1]$ $[1,1]$
@@
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-1,-0.988]$ $[0.988,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,-1]$ $[1,1]$
@@
Notes: See Section 4.1 for details about simulation exercise 2.
Table A.3: Estimated bounds under different joint covariate distributions, simulation exercise 3
Range of Lower Bounds Range of Upper Bounds
Parameter (1) (2)
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0,ELL_{i}=0\big{]}$ $[-0.222,-0.225]$ $[0.768,0.860]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0,ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=0\big{]}$ $[-1,-0.787]$ $[0.971,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=0\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=0$, $ELL_{i}=1\big{]}$ $[-0.897,-0.775]$ $[0.972,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=0$, $ELL_{i}=1\big{]}$
$\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=1,FRPL_{i}=1$, $ELL_{i}=1\big{]}$ $[-1,-0.796]$ $[0.959,1]$
$-\mathbb{E}\big{[}pass_{i}\big{|}white_{i}=0,FRPL_{i}=1$, $ELL_{i}=1\big{]}$
Notes: See Section 4.1 for details about simulation exercise 3.
We see that knowing the joint distribution of the covariates helps make estimated bounds (weakly) narrower for all parameters. Bounds on certain parameters can become a lot narrower when we know the joint distribution, like the first parameter in simulation exercise 1. While knowing the joint distribution of the covariates will not help in signing any of the marginal effect parameters, these results suggests that bounds do look different when joint covariate information is unavailable in the aggregate data.
Appendix B Appendix: Proofs
B.1 Proof of Lemma 1
@@ B.2 Proof of Proposition 1
$D=\bigg{\{}\sum_{k=1}^{K}\lambda_{k}d_{k}\bigg{|}~{}0\leq d_{k}\leq 1~{}\forall k,\text{ and }\exists\,(p_{1g},\dots,p_{Kg})\in P_{g},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }d_{k}=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}~{}\forall k,~{}\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g\bigg{\}},$ (B.1)
@@
$P_{g}=\bigg{\{}(p_{1g},\dots,p_{Kg})\in[0,1]^{K}\bigg{|}~{}\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{jg}~{}\forall\ell,k,\text{ and }\sum_{k=1}^{K}p_{kg}=1\bigg{\}}.$ (B.2)
If $\mathbb{P}[G_{i}=g],c_{kg}\in[0,1]$ for all $k$ and $g$ then because $d_{k}=\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}$ we know $d_{k}\in[0,1]$ for all $k$. Then we can get rid of $d_{k}$ in (B.1) by plugging in one of the constraints like so:
$D=\bigg{\{}\sum_{k=1}^{K}\lambda_{k}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]c_{kg}\bigg{|}~{}\exists\,(p_{1g},\dots,p_{Kg})\in P_{g},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g\bigg{\}}.$ (B.3)
@@
This means that we can equivalently express (B.3) and (B.2) as optimization problems as follows: $D=\left[L,U\right]$, where
@@
B.3 Proof of Proposition 2
Proof.
$\bigg{\{}\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}\bigg{|}~{}S_{g}c_{g}\leq a_{g}~{}\forall g\text{ and }\exists\,(p_{1g},\dots,p_{Kg})\in P_{g},(c_{1g},\dots,c_{Kg})\in[0,1]^{K}~{}\forall g\\ \text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g\bigg{\}}.$ (B.4)
We know that the set of $c_{g}\in[0,1]^{KG}$ such that $S_{g}c_{g}\leq a_{g}$ for all $g$ is a convex, closed, and bounded set because it is the intersection of a closed, bounded, and convex set with a half-space. This means the set of values of $\sum_{k}\lambda_{k}c_{kg}$ such that $S_{g}c_{g}\leq a_{g}$ for all $g$ is also a closed and bounded and convex set.
$\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}$ $\qquad\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}^{\prime}p_{kg}^{\prime}$ (B.5)
$S_{g}c_{g}\leq a_{g}$ $\qquad S_{g}c_{g}^{\prime}\leq a_{g},$ (B.6)
we know $tc_{kg}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$ and $tc_{kg}^{\prime}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$ also satisfy (B.5) in place of $c_{kg}$ and $c_{kg}^{\prime}$ respectively, for any $t\in[0,1]$. Thus for any $\sum_{k=1}^{K}\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}c_{kg}^{\prime}$, $\{\tilde{c}_{kg}\}$ satisfies $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}\tilde{c}_{kg}p_{kg}$. And that (B.6) holds implies $S_{g}(tc_{g}+(1-t)c_{g}^{\prime})\leq a_{g}$ for any $t\in[0,1]$, meaning that for any $\sum_{k=1}^{K}\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}c_{kg}^{\prime}$, $\{\tilde{c}_{kg}\}$ satisfies $S_{g}\tilde{c}_{g}\leq a_{g}$.
Thus any $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}^{\prime}$ is in the set given by (B.4), meaning (B.4) is an interval. Thus the sharp identified set is an interval. The rest of Proposition 2 can be proved following arguments similar to those used in the proof of Proposition 1. ∎
B.4 Proof of Proposition 3
Proof.
To prove sharpness of the identified set, it is sufficient to show that the following set is an interval:
$\bigg{\{}\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}\bigg{|}~{}\exists\,\{p_{kg}\}\in P_{g},\{c_{kg}\}\in[0,1]^{K}~{}\forall g\text{ s.t. }\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg}~{}\forall g,\\ \mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g},\forall g\bigg{\}}.$ (B.7)
Note that the set of $\{c_{kg}\}\in[0,1]^{KG}$ such that
$\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg}~{}\forall(\ell,k)\in F_{g},\forall g$ (B.8)
is a convex, closed, and bounded set because it is the intersection of a closed, bounded, and convex half-space. This means the set of values of $\sum_{k}\lambda_{k}c_{kg}$ such that (B.8) holds is also a closed and bounded and convex set.
In the proof of Proposition 1 I argued that for any $\{c_{kg}\},\{c_{kg}^{\prime}\}\in[0,1]^{K}$ with corresponding $\{p_{kg}\},\{p_{kg}^{\prime}\}\in P_{g}$ that satisfy
$\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}p_{kg},\qquad\qquad$ $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}c_{kg}^{\prime}p_{kg}^{\prime},$ (B.9)
$\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ $=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}p_{kg},$ (B.10)
$\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]$ $=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{kg}^{\prime}p_{kg}^{\prime},$ (B.11)
we know $tc_{kg}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$ and $tc_{kg}^{\prime}+(1-t)\mathbb{E}[Y_{i}|G_{i}=g]$ also satisfy (B.9) in place of $c_{kg}$ and $c_{kg}^{\prime}$ respectively, for any $t\in[0,1]$. Thus for any $\sum_{k=1}^{K}\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}c_{kg}^{\prime}$, $\{\tilde{c}_{kg}\}$ satisfies $\mathbb{E}[Y_{i}|G_{i}=g]=\sum_{k=1}^{K}\tilde{c}_{kg}p_{kg}$.
Note that, similar to the claim made in the proof of Proposition 1, since $\mathbb{P}[X_{\ell i}=x_{k,\ell}|G_{i}=g]=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{kg}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{kg}^{\prime}$ it follows that (B.10) still holds when letting $c_{kg}=\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$ and (B.11) still holds when letting $c_{kg}^{\prime}=\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$.
Thus $tc_{kg}+(1-t)\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$ in place of $c_{kg}$ and $tc_{kg}^{\prime}+(1-t)\mathbb{E}[Y_{i}|X_{\ell i}=x_{k,\ell},G_{i}=g]$ in place of $c_{kg}^{\prime}$ also satisfy (B.10) and (B.11) respectively for any $t\in[0,1]$. So for any $\sum_{k=1}^{K}\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\lambda_{k}c_{kg}^{\prime}$, $\{\tilde{c}_{kg}\}$ in place of $\{c_{kg}\}$ satisfies (B.10).
Thus any $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}\tilde{c}_{kg}$ between $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}$ and $\sum_{k=1}^{K}\sum_{g=1}^{G}\mathbb{P}[G_{i}=g]\lambda_{k}c_{kg}^{\prime}$ is in the set given by (B.7), meaning (B.7) is an interval. Thus the sharp identified set is an interval. The rest of Proposition 3 can be proved following arguments similar to those used in the proof of Proposition 1. ∎
B.5 Proof of Proposition 4
Proof.
I prove the result for $\hat{D}$ defined with respect to Proposition 1, then discuss the additional restrictions implied by Assumptions 2 and 3.
The result follows from the Theorem of the Maximum for the following correspondence:
$\Gamma\left(y,\{g_{k,\ell}\}_{k,\ell}\right)=\underset{\{p_{k},c_{k}\}_{k}}{\text{arg}\max}\left(\underset{\{p_{k},c_{k}\}_{k}}{\text{arg}\min}\right)\bigg{\{}\sum_{k=1}^{K}\lambda_{k}c_{k}\text{ s.t. }y=\sum_{k=1}^{K}c_{k}p_{k},\sum_{k=1}^{K}p_{k}=1,\\ \qquad g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j},c_{k}\in[0,1]~{}\forall k,p_{k}\geq 0~{}\forall k\bigg{\}}.$
Since $\sum_{k=1}^{K}p_{k}^{n}=1$ for all $n$, $p_{k}^{n}\geq 0$ for all $k,n$, and $p_{k}^{n}\to p_{k}$ for all $k$, it follows that $\sum_{k=1}^{K}p_{k}=1$ and $p_{k}\geq 0$ for all $k$. Since $c_{k}^{n}\in[0,1]$ for all $k,n$ and $c_{k}^{n}\to c_{k}$ for all $k$, it follows that $c_{k}\in[0,1]$ for all $k$. Since $y^{n}=\sum_{k=1}^{K}p_{k}^{n}c_{k}^{n}$ for all $k,n$ and $y^{n}\to y,d_{k}^{n}\to d_{k},p_{k}^{n}\to p_{k}$ for all $k,n$, it follows that $y=\sum_{k}p_{k}d_{k}$. Since $g_{k,\ell}^{n}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}^{n}$ for all $k,\ell,n$ and we have $p_{k}^{n}\to p_{k},g_{k,\ell}^{n}\to g_{k,\ell}$ for all $k,\ell$, it follows that $g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}$ for all $k,\ell$. Thus ($(\{p_{k}\},\{c_{k}\})\in\Gamma((y,\{g_{k,\ell}\}_{k,\ell}))$ and thus $\Gamma$ is upper hemicontinuous.
[…]
The law of large numbers also gives us that $\widehat{Pr}[G_{i}=g]$ is consistent, so that by continuous mapping theorem again we have that the lower and upper bounds of $\hat{D}$ converge to the lower and upper bounds of $D$.
To accommodate the restriction that $S_{g}c\leq a_{g}$ under Assumption 2 in correspondence $\Gamma$, note that any sequence $\{c_{k}^{n}\}\to\{c_{k}\}$ with ($(\{p_{k}^{n}\},\{c_{k}^{n}\})\in\Gamma\left((y^{n},\{g_{k,\ell}^{n}\}_{k,\ell})\right)$ satisfies $S_{g}c^{n}\leq a_{g}$ and thus $S_{g}c\leq a_{g}$ by continuity as well. So upper hemicontinuity is maintained.
The assumption that $\hat{D}$ is nonempty and the observed marginals are valid with respect to the assumed support ensures that for $(y^{n},\{g^{n}_{k,\ell}\}_{k,\ell})\to(y,\{g_{k,\ell}\}_{k,\ell})$ there exists ($(\{p_{k}\},\{c_{k}\})\in\Gamma(y,\{g_{k,\ell}\}_{k,\ell})$ and that $\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ is nonempty. Let ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ be a sequence such that $\max_{k}\{|p_{k}^{t}-p_{k}|,|c_{k}^{t}-c_{k}|\}$ is minimized over ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$ for each $t$.
I claim ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\to(\{p_{k}\},\{c_{k}\})$. From the restrictions that $g_{k,\ell}^{n_{t}}=\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}^{t}$ we know $\max_{k}\sum_{k=1}^{K}|p_{k}^{t}-p_{k}|\to 0$ as discussed above with the Moore-Penrose inverse solution to the linear system, so $\{p_{k}^{t}\}\to\{p_{k}\}$. And if $\max_{k}\sum_{k=1}^{K}|c_{k}^{t}-c_{k}|\not\to 0$ then $c_{k}^{t}p_{k}^{t}\not\to c_{k}p_{k}$, meaning the constraint $y=\sum_{k=1}^{K}c_{k}p_{k}$ cannot hold either as $y^{n_{t}}\to y$. Thus it must be that $\{c_{k}^{t}\}\to\{c_{k}\}$ and so lower hemicontinuity is also maintained because $S_{g}c^{t}\leq a_{g}$ by assumption.
To accommodate the restriction
$y_{k,\ell}g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{j}p_{j}$ (B.12)
under Assumption 3 in correspondence $\Gamma$, note that any sequence ($(\{p_{k}^{n}\},\{c_{k}^{n}\})\to(\{p_{k}\},\{c_{k}\})$ with ($(\{p_{k}^{n}\},\{c_{k}^{n}\})\in\Gamma\left((y^{n},\{g_{k,\ell}^{n}\}_{k,\ell},\{y_{k,\ell}^{n}\}_{k,\ell})\right)$ satisfies
$y_{k,\ell}^{n}g_{k,\ell}^{n}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{j}^{n}p_{j}^{n},$
and thus (B.12) holds by continuity as well. So upper hemicontinuity is maintained.
The assumption that $\hat{D}$ is nonempty and the observed marginals are valid with respect to the assumed support ensures that for $(y^{n},\{g^{n}_{k,\ell}\}_{k,\ell},\{y^{n}_{k,\ell}\}_{k,\ell})\to(y,\{g_{k,\ell}\}_{k,\ell},\{y_{k,\ell}\}_{k,\ell})$ there exists ($(\{p_{k}\},\{c_{k}\})\in\Gamma(y,\{g_{k,\ell}\}_{k,\ell},\{y_{k,\ell}\}_{k,\ell})$ and that $\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell},\{y_{k,\ell}^{n_{t}}\}_{k,\ell})$ is nonempty. Let ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell},\{y_{k,\ell}^{n_{t}}\}_{k,\ell})$ be a sequence such that $\max_{k}\{|p_{k}^{t}-p_{k}|,|c_{k}^{t}-c_{k}|\}$ is minimized over ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\in\Gamma(y^{n_{t}},\{g_{k,\ell}^{n_{t}}\}_{k,\ell})$for each $t$.
I claim ($(\{p_{k}^{t}\},\{c_{k}^{t}\})\to(\{p_{k}\},\{c_{k}\})$. From the restrictions that $g_{k,\ell}^{n_{t}}=\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}p_{j}^{t}$ we know $\max_{k}\sum_{k=1}^{K}|p_{k}^{t}-p_{k}|\to 0$ as discussed above with the Moore-Penrose inverse solution to the linear system. And if $\max_{k}\sum_{k=1}^{K}|c_{k}^{t}-c_{k}|\not\to 0$ then $c_{k}^{t}p_{k}^{t}\not\to c_{k}p_{k}$, meaning the constraints $y=\sum_{k=1}^{K}c_{k}p_{k}$ and $y_{k,\ell}g_{k,\ell}=\sum_{j=1}^{K}\mathbbm{1}\{x_{j,\ell}=x_{k,\ell}\}c_{j}p_{j}$ cannot hold either as $y^{n_{t}}\to y,g_{k,\ell}^{n_{t}}\to g_{k,\ell},y_{k,\ell}^{n_{t}}\to y_{k,\ell}$. Thus it must be that $\{c_{k}^{t}\}\to\{c_{k}\}$ and so lower hemicontinuity is also maintained. ∎
B.6 Proof of Proposition 5
Proof.
B.7 Proof of Corollary A.1